OpenStudy (hugy0212):

Convert y=4x^2-4 to vertex form

OpenStudy (anonymous):

It's already in vertex form

OpenStudy (jdoe0001):

\( y=4x^2-4 \implies y=4(x-{\color{brown}{ 0}})^2{\color{blue}{ -4}} \\ \quad \\ y=(x-{\color{brown}{ h}})^2+{\color{blue}{ k}}\\ x=(y-{\color{blue}{ k}})^2+{\color{brown}{ h}}\qquad\qquad vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\)